=Yesterday, Manuel's electronics store sold several radios at $18.50 each and a few blank cassettes at $1 each. The difference in the number of these items sold was eight. If the total sales of radios and cassettes for the day were $265.00, how many of each item were sold? Use the addition method of solving systems.
Answer by checkley75(3666)

1850x+100y=26500
x-y=8 multiply this equation by 100 & add.
100x-100y=800
1850x+100y=26500
------------------------------------------------
1950x=27300
x=27300/1950
x=14 radios were sold.
therefore there were 14-y=8 or -y=8-14 or y=6 cassettes sold.
proof
14*1850+6*100=26500
25900+600=26500
26500=26500